archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    Incident Category Status CLOSED Percent Complete Details The vac2 Arbor equipment failed vac2 was cut traffic is redirected on vac1 Date Thursday 29 January 2015 19 13PM Reason for closing Done Comments 1 Comment by OVH Thursday 29 January 2015

    Original URL path: http://status.ovh.es/?do=details&id=8647 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    CLOSED Percent Complete Details We are improving network firewall activation No cut in service is expected for IPs with network firewall activated Date Tuesday 27 January 2015 02 00AM Reason for closing Done Comments 1 Comment by OVH Tuesday 27

    Original URL path: http://status.ovh.es/?do=details&id=8615 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Details To better debug the task 12258 http status ovh net do details id 8387 we suspended the operation on the ON and OFF for the permanent mitigation In 48 hours we are going to put them back During the 48 hours you won t be able to add or remove ips on the permanent mitigation If the problem comes from there we are going to redo the ON and

    Original URL path: http://status.ovh.es/?do=details&id=8435 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Let s go We are first disabling the upgrade feature on the site as well as the creation of new accounts Then we will stop connections the time of the data update 3 to 4 hours The oAuth protocol will be activated in production We will then install the new version of the web application A new connection password will be requested from you to be able to use the new API oAuth On Friday it will be possible to create and upgrade accounts Comment by OVH Friday 04 October 2013 11 15AM All connections to web services are now stopped The new connections in applications will be stopped Comment by OVH Friday 04 October 2013 11 16AM The web services are now stopped For the persons who are already connected all remains functional Comment by OVH Friday 04 October 2013 11 16AM All robots are stopped account creation upgrade and publishing Comment by OVH Friday 04 October 2013 11 18AM We are now preparing the database to receive all accounts The operation is in 3 steps Retrieve data from the current database Formate data to insert it in the new infrastructure Retrieve data from the new infrastructure Comment by OVH Friday 04 October 2013 11 18AM Step 1 done Step 2 doing Comment by OVH Friday 04 October 2013 11 18AM Step 2 done Step 3 doing Comment by OVH Friday 04 October 2013 11 19AM 75 of the database is imported Comment by OVH Friday 04 October 2013 11 20AM Import is done We are now testing it before reactivating the web services The aim is to keep the compatibility with the current applications Comment by OVH Friday 04 October 2013 11 20AM Tests are positive Within 1 hour the web services will be available again Comment by

    Original URL path: http://status.ovh.es/?do=details&id=5525 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    FS 13111 filerz2006 Attached to Project Hosting Task Type Modernization Category 1000gp Status CLOSED Percent Complete Details There are network errors at the card exit We will investigate Date Monday 20 April 2015 18 35PM Reason for closing Done Comments

    Original URL path: http://status.ovh.es/?do=details&id=9252 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Incident Category 1000gp Status CLOSED Percent Complete Details The DAS disks are no longer being detected and intervention is underway Date Monday 01 December 2014 09 50AM Reason for closing Done Additional comments about closing An expander was the cause

    Original URL path: http://status.ovh.es/?do=details&id=8251 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Incident Category 1000gp Status CLOSED Percent Complete Details Our monitoring system has detected instabilities They were due to a sensor that panicked and unintentionally rebooted the Apache httpd server We ve corrected the sensor cluster012 ois stabilised Date Monday 22

    Original URL path: http://status.ovh.es/?do=details&id=7754 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    the host server the navigator displays the webmail Very soon a page will display It will highlight that the domain has not yet been declared on the host server helping you therefore to carry out a diagnostic on and correct your website We will begin with cluster012 1000gp Date Tuesday 30 December 2014 17 20PM Reason for closing Done Comments 3 Comment by OVH Tuesday 23 September 2014 13 30PM

    Original URL path: http://status.ovh.es/?do=details&id=7733 (2015-10-15)
    Open archived version from archive



  •